There were 1500 people at a high school football game. Student tickets were $2.00 and adult tickets were $3.50. The total receipts for the game were $3825. How many adult and student tickets were sold?
Hint: Again two equations for two unknowns. Write one for the number of tickets and one for the value of the tickets. You're on your own this time.

Answer by ptaylor(1894)


Let x=number of adult tickets
Then 1500-x=number of student tickets
Now we are told that the receipts taken in from selling adult tickets ($3.50x)
plus the receipts taken in from selling student tickets [$2.00(1500-x)] equals $3825. So our equation to solve is:
$3.50x+$2.00(1500-x)=$3825
$3.50x+$3000-$2.00x=$3825 subtract $3000 from both sides:
$1.50x=$3825-$3000
$1.50x=$825
x=550 adult tickets sold
1500-x=1500-550=950 student tickets sold
Ck
550($3.50)+950($2.00)=$3825
$1925+$1900=$3825
$3825=$3825



Hope this helps-----ptaylor